# The Generic Temperature Response of Large Biochemical Networks

### v2

@@ I Introduction
Because developmental systems are often shielded from temperature changes (e.g. due to an egg shell or a placenta which protects the embryo), they seem to have evolved less temperature compensation and therefore appear to be better model systems to study the effect of temperature changes on large biochemical networks [47, 48]. Measurements of the prepupal development time of Drosophila under different temperature conditions were published almost one hundred years ago by Bliss [25]. The results clearly showed a concave curve in the Arrhenius plot, as opposed to the straight line usually found for single reactions. About a decade later, Powsner [26] obtained similar results for the embryonic, larval and pupal phases of the development of Drosophila, already hypothesizing that this may be a consequence of the complexity of the underlying network. More recently, Crapse et al. [24] published a study in which the embryonic development of Drosophila was subdivided into twelve phases. The authors measured the temperature dependence of the completion time of each of them. On this level, a concave temperature response emerged as well, which could be accurately described by a quadratic fit in the Arrhenius plot. In the same study, similar results were also found for Xenopus laevis embryos. The authors also performed simulations for a linear sequence of Poisson processes with statistically distributed parameters, but the results could not fully reproduce the shapes of the experimental curves [24]. However, this approach disregards the fact that the underlying biochemical networks are bound to be highly non-linear, with complicated feedback mechanisms and network motifs [49]. In general, although there has been some work on understanding the temperature dependence of development rates from first principles [50, 51, 52, 53, 54], the main body of work has remained empirical [55, 56, 57, 58, 59, 60].
@@
The general question addressed in this work is illustrated in Fig. 1. The temperature dependence of a single reaction is usually described well by the Arrhenius equation, resulting in a linear relationship in the Arrhenius plot (Fig. 1a). However, when considering complex biological processes, e.g. the whole mitogen-activated protein kinase (MAPK) signaling pathway triggered by epidermal growth factors (EGF), that transmits extracellular signals to the nucleus and regulates differentiation [61], the linear relation is lost and microscopic details might govern the response (Fig. 1b). In the limit of very large systems, like whole organisms such as Drosophila developing from a fertilized egg to a larva [62], the overall rate usually approaches a quadratic dependence (Fig. 1c). Here we provide a rigorous mathematical derivation of why large biochemical networks typically show a quadratic shape in the Arrhenius plot. Our proof is based on calculating first passage times using a graph-theoretical decomposition of the network into spanning trees and forests. Our result is valid for all network architectures, including feedback loops, as long as several spanning trees stay relevant in the limit of large system size. Thus, our work solves the long-standing question of what the generic temperature response of large and complicated biochemical networks is. To support our analytical findings, we have developed a numerical procedure that generates networks of given sizes and with a bias towards an endpoint, and then uses the graph-theoretical interpretation of mean first-passage times to generate their distribution. This numerical procedure confirms the generic temperature response predicted by our theory. We also show that our results are in excellent agreement with experimental data from developmental systems like fly and frog. Finally we discuss the cases in which the generic response does not occur, including linear chains, which have only one spanning tree. Our approach paves the way to deal with smaller biochemical systems and more specific temperature responses, including protein denaturation.
@@ II.1 Meaning of mean first-passage time
A general framework to mathematically describe a chemical reaction network is to express it as a time-dependent probability distribution $p_{i}(t)$ on $N_{v}$ discrete states, where without loss of generality we denote states $1$ and $N_{v}$ as start and target states, respectively. For a linear chain, we would have $N=N_{v}-1$ reactions, each with forward and backward directions, while in general, the system could contain many loops. For our theoretical considerations, one needs both $N+1$ and $N$, thus we first use $N$ as main variable; for the numerical part coming later, it is more convenient to use $N_{v}=N+1$.
@@ II.2.2 Graph-theoretical interpretation of FPT moments
where $\mathcal{T}_{[N+1]}$ denotes the spanning trees of the corresponding graph rooted at $N+1$ and their weights are defined as $w(\mathcal{T}):=\prod_{k_{ij}\in E(\mathcal{T})}k_{ij}$, with $E(\mathcal{T})$ being the edge set of $\mathcal{T}$. Fig. (3) shows $\mathcal{T}_{[4]}$ for the complete graph on four vertices as an example.
@@
where $\mathcal{F}_{[j,N+1]}^{i\rightarrow j}$ are the spanning forests with two trees of the graph, one rooted at $j$ and containing $i$ and the other one rooted at $N+1$. Fig. (4) shows $\mathcal{F}_{[i,4]}^{1\rightarrow i}$ for the complete graph on 4 vertices.
@@ II.3 Effect of temperature
Since most biological organisms have evolutionarily adapted to a relatively fixed thermal environment, it is reasonable to rephrase the Arrhenius equation in terms of the rate constant at a reference temperature $T_{0}$, introducing $k^{0}_{i}=k_{i}(T=T_{0})$. $T_{0}$ describes the temperature at which the organism usually operates. For instance, for endothermic animals, the body temperature would serve as the natural reference point, while for exothermic animals, it would be the typical temperature of the environment. The $k^{0}_{i}$ can be seen as the standard rates, and a crucial assumption for the following analysis is that these standard rates are independent of the activation energies $E_{i}$.
The Arrhenius equation can be expressed in terms of these parameters as:
$k_{ij}$ $=A_{ij}e^{-\frac{E_{ij}}{k_{B}T}}=\underbrace{k_{ij}(T=T_{0})}_{=:k^{0}_{ij}}e^{-\frac{E_{ij}}{k_{B}}(\frac{1}{T}-\frac{1}{T_{0}})},$ (22)
@@ II.4 Partition sums and generating functions
Letting the sums in Eq. (23) and Eq. (24) run only over the non-vanishing trees, all involved rates satisfy $k_{ij}>0$, and one can parameterize them as $k_{ij}(a)=e^{aX_{i,j}}$ such that $k_{ij}(a=1)=e^{X_{i,j}}$ is the rate of interest. Applying this parametrization, one obtains for $Z_{\mathcal{T}}$:
@@
defining $X_{\mathcal{T}}:=\sum_{k_{ij}\in E(\mathcal{T})}X_{ij}$, which is the sum over all $X_{ij}$ of a given tree. Now consider the Taylor-series of $\ln{Z_{\mathcal{T}}}$ around $a=0$:
$\ln{Z_{\mathcal{T}}}(a)=\sum_{n=0}^{\infty}\frac{\partial^{n}\ln{Z_{\mathcal{T}}}}{\partial^{n}a}\Big{|}_{a=0}\frac{a^{n}}{n!}$ (27)
@@ II.5 Limit of large system size
Note that backward rates do not fully enter in the spanning trees (cf. examples in Appendix A) so the activation energies on the trees and on the two-tree forests are different in general. The result so far requires to know the cumulants $\kappa_{n}^{E_{\mathcal{T}}}$ and $\kappa_{n}^{E_{\mathcal{F}}}$ to get the global rate constant.
@@
$E_{\mathcal{T}}=\sum_{k_{ij}\in E(\mathcal{T})}E_{ij}$ and $E_{\mathcal{F}}=\sum_{k_{ij}\in E(\mathcal{F})}E_{ij}$ are the sums of $N$, and N1N-1 independent random variables and, since $N\gg 1$, the central limit theorem suggests indeed that this is a valid assumption as long as the higher cumulants vanish fast enough compared to the differences of the first two terms, requiring sufficiently well-behaved convergence to the normal distribution. The simplest case is that the $E_{ij}$ already follow a normal distribution, which may be different for the $E_{ij}$ on the trees and the two tree-forests, their sums also follow a normal distribution and one obtains:
@@ II.6 Comparison to simulated networks
To illustrate the convergence predicted by theory, random networks on $N_{v}=10$ to $N_{v}=1.000$ vertices displaying some basic characteristics of the biochemical network of a developmental process were generated (see Appendix D for details) [83]. In particular, they were designed to have a relatively large distance between the initial vertex 0 and the final vertex $N_{v}$ as well as a block-like structures with many connections and irreversible transitions between them, reflecting the checkpoints present in development. Two small networks with $N_{v}=30$ and $N_{v}=50$ are shown as examples in Fig. 5a and b, respectively.
@@
The spanning trees and two-tree spanning forests of these base networks were then drawn at random. A standard procedure to generate random spanning trees is given by Wilson’s algorithm [84]. While removing a random edge from a spanning tree will result in a spanning forest, this procedure does not yield all spanning forests with the same probability. In order to ensure an unbiased sample, a specific algorithm was devised for the generation of two-tree spanning forests. In Appendix C, the steps needed for this algorithm are explained and its functionality is illustrated for a simple example network.
@@
In each case, $10.000$ spanning trees and at least $10.000$ two-tree spanning forests (note that the number of generated forests is not predefined to ensure an unbiased sample, as explained in Appendix C), and the total activation energies for each of them were calculated. The total activation energies for $N_{v}=10$ are shown in Fig. 6 as histograms. One observes that there are just a few spanning trees and also a moderate number of two-tree spanning forests, so that the distribution is not reminiscent of a Gaussian yet. The other histograms, along with fits to normal distributions, are shown in Fig. 7. As expected, the histograms of the total activation energies approach a normal distribution as $N_{v}$ increases.
Then, Eq. (37) can be applied to obtain the Arrhenius plot for the global rate. The corresponding plots for networks of size $20$, $100$ and $1.000$ are shown in Fig. 7. The convergence of the total energy distributions to a Gaussian is observed as the system size $N_{v}$ increases. For the parameters chosen here, the distribution appears already well described by a Gaussian for $N_{v}\sim 100$. The second derivative, i.e., twice the quadratic term as a function of $N_{v}$ is shown in Fig. (8). One observes that as the system size increases, so does the quadratic term, in a statistical manner. This proves numerically that this kind of networks generically lead to a quadratic Arrhenius plot in the limit of large system size.
@@
Figure 9: Arrhenius plot for the rates to the different stages in the development of Drosophila melanogaster embryos, starting at the 14th nuclear division (A) adapted from [24] with quadratic fit to data. The residuals to the fit are shown below the main diagram. The labeling of the stages follows the convention from the original authors: (a) rate from A to B (beginning of cellularization), $R^{2}=0.9587$ (b) rate from A to E (development of horizontal posterior midgut), $R^{2}=0.9909$ (c) rate from A to H (full germ band retraction), $R^{2}=0.9952$ (d) rate from A to K (first breath), $R^{2}=0.9868$.
Figure 10: Logarithmic developmental rate to the pupal stage of Drosophila melanogaster versus inverse temperature adapted from [25] with quadratic fit. The residuals to the fit are shown below the main diagram. (a) For male specimen, $R^{2}=0.9950$ (b) for female specimen, $R^{2}=0.9950$.
@@ II.7 Comparison to experimental data
Our theoretical prediction is in striking agreement with experimental data for developmental systems. In a work published in 2021, Crapse et al. [24] conducted a study on the development rates of Drosophila embryos at different temperatures. To that end, the embryonic phase of Drosophila was subdivided into 12 stages and the rates to reach each stage were measured under different temperature conditions. The 14th nuclear division in the zygote was chosen as the starting point (stage A) and the first breath was chosen as the end point (stage K). The cumulative rates starting with the 14th nuclear division to 4 different stages (B,E,H,K in the notation of the original authors) are shown in Fig. 9 as an example, together with a quadratic fit performed in Python via the numpy.polyfit command. The data were taken from the supplementary material of the original publication, which tabulated the measured times for the entire sample and all stages. The fitting parameters are listed in Appendix B. Note that these do not provide insights into the microscopic structure of the network without further assumptions since the coefficients of the quadratic functions contain the differences of the first two moments along the trees and forests as well as terms containing $\frac{1}{k_{B}T_{0}}$. Moreover, the residuals of the quadratic fit were computed to analyze whether there are any systematic deviations that hint towards a deviation from the quadratic dependence.
@@
Fig. 12 shows the quadratic terms of the quadratic fit for the rates to the different stages defined by Crapse et al. Consistent with the simulation data shown in Fig. 8, the quadratic term tends to increase as more steps accumulate. However, a slight decrease is observed for the rates to the final stages. This is most likely due to the lack of data points at the extreme ends of the temperature spectrum, as most embryos did not survive long enough under these conditions. Consequently, this also results in greater uncertainties for these values.
Measurements of the temperature dependence of developmental rates of Drosophila were already performed by Bliss [25] in 1926 for the prepupal development and by Powsner [26] in 1935 for the embryonic, pupal, and larval stages. Their results are shown in Fig. 10 and 11, respectively, also with a quadratic fit to their data points. Both articles list their data in a table and show them in a diagram, albeit together with piecewise linear functions as fit instead of a global quadratic function. One sees that in all cases, the quadratic fits predicted by our theory are in very good agreement with the experimental data. Also the residuals indicate only a minor contributions from terms of higher order, except for the larval development rates from the data by Powsner (Fig. 11 (b,c)), where most likely some additional effect contributes.
Moreover, one also observes that a clear quadratic shape becomes more prominent the more steps accumulate, as also anticipated from the previous theoretical considerations. Note that the maximal rate, i.e. the peak in the diagrams, is at $T_{peak}\sim 30^{\circ}\text{C}$. Expecting the mean total activation energies to be higher for the trees than for the forests, and vice-versa for the variances, Eq. (37) predicts the maximum to be attained at $\Delta\beta<0$, so $T_{peak}>T_{0}$, and one would in fact assume $T_{0}\sim 20^{\circ}\text{C}$.

### v3

@@ I Introduction
Developmental systems seem to have evolved less temperature compensation and therefore appear to be better model systems to study the effect of temperature changes on large biochemical networks [47, 48]. Measurements of the prepupal development time of Drosophila under different temperature conditions were published almost one hundred years ago by Bliss [25]. The results clearly showed a concave curve in the Arrhenius plot, as opposed to the straight line usually found for single reactions. About a decade later, Powsner obtained similar results for the embryonic, larval and pupal phases of the development of Drosophila, already hypothesizing that this may be a consequence of the complexity of the underlying network [26]. More recently, Crapse et al. published a study in which the embryonic development of Drosophila was subdivided into twelve phases [24]. The authors measured the temperature dependence of the completion time of each of them. On this level, a concave temperature response emerged as well, which could be accurately described by a quadratic fit in the Arrhenius plot. In the same study, similar results were also found for Xenopus laevis embryos. The authors also performed simulations for a linear sequence of Poisson processes with statistically distributed parameters, but the results could not fully reproduce the shapes of the experimental curves [24]. However, this approach disregards the fact that the underlying biochemical networks are bound to be highly non-linear, with complicated feedback mechanisms and network motifs [49]. Even more recently, Rombouts et al. confirmed the quadratic shape in the Arrhenius plot of the development rates of Xenopus embryos and also observed it for the development of zebra fish (Danio rerio) embryos, while for Caenorhabditis elegans the rate shows a more biphasic behavior [48]. In general, although there has been some work on understanding the temperature dependence of development rates from first principles [50, 51, 52, 53, 54], the main body of work has remained empirical [55, 56, 57, 58, 59, 60].
@@
The general question addressed in this work is illustrated in Fig. 1. The temperature dependence of a single reaction is usually described well by the Arrhenius equation, resulting in a linear relationship in the Arrhenius plot (Fig. 1a). However, when considering complex biological processes, e.g. the whole mitogen-activated protein kinase (MAPK) signaling pathway triggered by epidermal growth factors (EGF), that transmits extracellular signals to the nucleus and regulates differentiation [61], the linear relation is lost and microscopic details might govern the response (Fig. 1b). In the limit of very large systems, like whole organisms such as Drosophila developing from a fertilized egg to a larva [62], the overall rate usually approaches a quadratic dependence (Fig. 1c). Here we provide a rigorous mathematical derivation of why large biochemical networks typically show a quadratic shape in the Arrhenius plot. Our proof is based on calculating first-passage times using a graph-theoretical decomposition of the network into spanning trees and forests. Our result is valid for all network architectures, including feedback loops, as long as several spanning trees stay relevant in the limit of large system size. Thus, our work solves the long-standing question of what the generic temperature response of large and complicated biochemical networks is. To support our analytical findings, we have developed a numerical procedure that generates networks of given sizes and with a bias towards an endpoint, and then uses the graph-theoretical interpretation of mean first-passage times to generate their distribution. This numerical procedure confirms the generic temperature response predicted by our theory. We also show that our results are in excellent agreement with experimental data from developmental systems like fly and frog. Finally we discuss the cases in which the generic response does not occur, including linear chains, which have only one spanning tree. Our approach paves the way to deal with smaller biochemical systems and more specific temperature responses, including protein denaturation.
@@ II.1 Meaning of mean first-passage time
A general framework to mathematically describe a chemical reaction network is to express it as a time-dependent probability distribution $p_{i}(t)$ on $N_{v}$ discrete states, where without loss of generality we denote states $1$ and $N_{v}$ as start and target states, respectively. For a linear chain, we would have $N=N_{v}-1$ reactions, each with forward and backward directions, while in general, the system could contain many loops. In the following, we will first use $N$ as main variable; for the numerical part coming later, it is more convenient to use $N_{v}=N+1$.
@@ II.2.2 Graph-theoretical interpretation of FPT moments
where $\mathcal{T}_{[N+1]}$ denotes the spanning trees of the corresponding graph rooted at $N+1$ and their weights are defined as $w(\mathcal{T}):=\prod_{k_{ij}\in E(\mathcal{T})}k_{ij}$, with $E(\mathcal{T})$ being the edge set of $\mathcal{T}$. Fig. 3 shows $\mathcal{T}_{[4]}$ for the complete graph on four vertices as an example.
@@
where $\mathcal{F}_{[j,N+1]}^{i\rightarrow j}$ are the spanning forests with two trees of the graph, one rooted at $j$ and containing $i$ and the other one rooted at $N+1$. Fig. 4 shows $\mathcal{F}_{[i,4]}^{1\rightarrow i}$ for the complete graph on 4 vertices.
@@ II.3 Effect of temperature
Since most biological organisms have evolutionarily adapted to a relatively fixed thermal environment, it is reasonable to rephrase the Arrhenius equation in terms of the rate constant at a reference temperature $T_{0}$, introducing $k^{0}_{ij}=k_{ij}(T=T_{0})$. $T_{0}$ describes the temperature at which the organism usually operates. For instance, for endothermic animals, the body temperature would serve as the natural reference point, while for exothermic animals, it would be the typical temperature of the environment. The $k^{0}_{ij}$ can be seen as the standard rates, and a crucial assumption for the following analysis is that these standard rates are independent of the activation energies $E_{ij}$. The Arrhenius equation can then be expressed in terms of these parameters as:
$k_{ij}$ $=A_{ij}e^{-\frac{E_{ij}}{k_{B}T}}=k^{0}_{ij}e^{-\Delta\beta E_{ij}}$ (22)
@@ II.4 Partition sums and generating functions
Letting the sums in Eq. (23) and Eq. (24) run only over the non-vanishing trees, all involved rates satisfy $k_{ij}>0$, and one can parameterize them as $k_{ij}(a)=e^{aX_{ij}}$ such that $k_{ij}(a=1)=e^{X_{ij}}$ is the rate of interest. Applying this parametrization, one obtains for $Z_{\mathcal{T}}$:
@@
defining $X_{\mathcal{T}}:=\sum_{k_{ij}\in E(\mathcal{T})}X_{ij}$, which is the sum over all $X_{ij}$ of a given tree. Now consider the Taylor series of $\ln{Z_{\mathcal{T}}}$ around $a=0$:
$\ln{Z_{\mathcal{T}}}(a)=\sum_{n=0}^{\infty}\frac{\partial^{n}\ln{Z_{\mathcal{T}}}}{\partial^{n}a}\Big|_{a=0}\frac{a^{n}}{n!}\ .$ (27)
@@ II.5 Limit of large system size
Note that backward rates do not fully enter in the spanning trees (cf. examples in Appendix A), so the activation energies on the trees and on the two-tree forests are different in general. The result so far requires to know the cumulants $\kappa_{n}^{E_{\mathcal{T}}}$ and $\kappa_{n}^{E_{\mathcal{F}}}$ to get the global rate constant.
@@
$E_{\mathcal{T}}=\sum_{k_{ij}\in E(\mathcal{T})}E_{ij}$ and $E_{\mathcal{F}}=\sum_{k_{ij}\in E(\mathcal{F})}E_{ij}$ are the sums of $N$ and N1N-1 independent random variables, respectively. Since $N\gg 1$, the central limit theorem suggests that this is a valid assumption as long as the higher cumulants vanish fast enough compared to the differences of the first two terms, requiring sufficiently well-behaved convergence to the normal distribution. The simplest case is that the $E_{ij}$ already follow a normal distribution. However, any distribution, including the uniform one, is sufficient, as long as it satisfies the conditions required for the central limit theorem. Then one obtains:
@@ II.6 Comparison to simulated networks
To illustrate the convergence predicted by our theory, random networks on $N_{v}=10$ to $N_{v}=1.000$ vertices displaying some basic characteristics of the biochemical network of a developmental process were generated (see Appendix C for details) [83]. In particular, they were designed to have a relatively large distance between the initial vertex 0 and the final vertex $N_{v}=N+1$ as well as a block-like structures with many connections and irreversible transitions between them, reflecting the checkpoints present in development. Two small networks with $N_{v}=30$ and $N_{v}=50$ are shown as examples in Fig. 5a and b, respectively.
@@
The spanning trees and two-tree spanning forests of these base networks were then drawn at random. A standard procedure to generate random spanning trees is given by Wilson’s algorithm [84]. While removing a random edge from a spanning tree will result in a spanning forest, this procedure does not yield all spanning forests with the same probability. In order to ensure an unbiased sample, a specific algorithm was devised for the generation of two-tree spanning forests. In Appendix D, the steps needed for this algorithm are explained and its functionality is illustrated for a simple example network.
@@
In each case, $10.000$ spanning trees and at least $10.000$ two-tree spanning forests (note that the number of generated forests is not predefined to ensure an unbiased sample, as explained in Appendix D) and the total activation energies for each of them were calculated. The total activation energies for $N_{v}=10$ are shown in Fig. 6 as histograms. One observes that there are just a few spanning trees and also a moderate number of two-tree spanning forests, so that the distribution is not reminiscent of a Gaussian yet. The other histograms, along with fits to normal distributions, are shown in Fig. 7. As expected, the histograms of the total activation energies approach a normal distribution as $N_{v}$ increases.
Then, Eq. (37) can be applied to obtain the Arrhenius plot for the global rate. The corresponding plots for networks of size $20$, $100$ and $1.000$ are shown in Fig. 7. The convergence of the total energy distributions to a Gaussian is observed as the system size $N_{v}$ increases. For the parameters chosen here, the distribution appears already well described by a Gaussian for $N_{v}\sim 100$. The second derivative, i.e., twice the quadratic term as a function of $N_{v}$ is shown in Fig. 8. One observes that as the system size increases, so does the quadratic term, in a statistical manner. This proves numerically that this kind of networks generically lead to a quadratic Arrhenius plot in the limit of large system size.
@@
Figure 9: Arrhenius plot for the rates to the different stages in the development of Drosophila melanogaster embryos, starting at the 14th nuclear division (A) adapted from [24] with quadratic fit to data. The residuals to the fit are shown below the main diagram.
The labeling of the stages follows the convention from the original authors: (a) rate from A to B (beginning of cellularization), $R^{2}=0.9587$ (b) rate from A to E (development of horizontal posterior midgut), $R^{2}=0.9909$ (c) rate from A to H (full germ band retraction), $R^{2}=0.9952$ (d) rate from A to K (first breath), $R^{2}=0.9868$.
Figure 10: Logarithmic developmental rate to the pupal stage of Drosophila melanogaster versus inverse temperature adapted from [25] with quadratic fit. The residuals to the fit are shown below the main diagram.
(a) For male specimen, $R^{2}=0.9950$ (b) for female specimen, $R^{2}=0.9950$.
@@ II.7 Comparison to experimental data
Our theoretical prediction is in striking agreement with experimental data for developmental systems. In a work published in 2021, Crapse et al. conducted a study on the development rates of Drosophila embryos at different temperatures [24]. To that end, the embryonic phase of Drosophila was subdivided into 12 stages and the rates to reach each stage were measured under different temperature conditions. The 14th nuclear division in the zygote was chosen as the starting point (stage A) and the first breath was chosen as the end point (stage K). The cumulative rates starting with the 14th nuclear division to 4 different stages (B,E,H,K in the notation of the original authors) are shown in Fig. 9 as an example, together with a quadratic fit performed in Python via the numpy.polyfit command. The data were taken from the supplementary material of the original publication, which tabulated the measured times for the entire sample and all stages. The fitting parameters are listed in Appendix B. Note that these do not provide insights into the microscopic structure of the network without further assumptions, since the coefficients of the quadratic functions contain the differences of the first two moments along the trees and forests as well as terms containing $\frac{1}{k_{B}T_{0}}$. Moreover, the residuals of the quadratic fit were computed to analyze whether there are any systematic deviations that hint towards a deviation from the quadratic dependence.
@@
Measurements of the temperature dependence of developmental rates of Drosophila were already performed by Bliss [25] in 1926 for the prepupal development and by Powsner [26] in 1935 for the embryonic, pupal, and larval stages. Their results are shown in Fig. 10 and 11, respectively, also with a quadratic fit to their data points. Both articles list their data in a table and show them in a diagram, albeit together with piecewise linear functions as fit instead of a global quadratic function. One sees that in all cases, the quadratic fits predicted by our theory are in excellent agreement with the experimental data. Also the residuals indicate only a minor contributions from terms of higher order, except for the larval development rates from the data by Powsner (Fig. 11b,c), where most likely some additional effect contributes.
Fig. 12 shows the quadratic terms of the quadratic fit for the rates to the different stages defined by Crapse et al. [24]. Consistent with the simulation data shown in Fig. 8, the quadratic term tends to increase as more steps accumulate. However, a slight decrease is observed for the rates to the final stages. This is most likely due to the lack of data points at the extreme ends of the temperature spectrum, as most embryos did not survive long enough under these conditions. Consequently, this also results in greater uncertainties for these values.
Note that the maximal rate, i.e. the peak in the diagrams, is at $T_{peak}\sim 30^{\circ}\text{C}$. Expecting the mean total activation energies to be higher for the trees than for the forests, and vice-versa for the variances, Eq. (37) predicts the maximum to be attained at $\Delta\beta<0$, so $T_{peak}>T_{0}$, and one would in fact assume $T_{0}\sim 20^{\circ}\text{C}$.